# Noise estimation of SDE from a single data trajectory

### v1

@@ 3 Main Results
$+\psi_{21}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ dB^{\mathbb{Q}}_{r}du+\psi_{22}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ dB^{\mathbb{Q}}_{r}\circ dB_{u}^{\mathbb{Q}}+\psi_{222}^{s,t}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}+\mathsf{\varepsilon}^{\Psi_{2}},$
@@
3.2 Proof of Step 3.
Proof: Applying stochastic Taylor expansion on $\tilde{\mu}$ and $\sigma$ around $X_{s}$, and truncating the Taylor expansion at level 2, we have the following approximation of $\tilde{\mu}$ and $\sigma$,
$\tilde{\mu}(X_{u})$ $=\tilde{\mu}(X_{s})+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dX_{r}+\underbrace{\int_{s}^{u}\int_{s}^{r}\int_{s}^{v}\tilde{\mu}^{\prime\prime\prime}(\xi_{\tau})\circ dX_{\tau}\circ dX_{v}\circ dX_{r}}_{\varepsilon^{\tilde{\mu}}}$ (30)
$\sigma(X_{u})$ $=\sigma(X_{s})+\sigma^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\sigma^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dX_{r}+\underbrace{\int_{s}^{u}\int_{s}^{r}\int_{s}^{v}\sigma^{\prime\prime\prime}(\eta_{\tau})\circ dX_{\tau}\circ dX_{v}\circ dX_{r}}_{\varepsilon^{\sigma}}$ (31)
@@
$X_{t}$ $=X_{s}+\int_{s}^{t}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{s})+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}{\circ}dX_{v}\circ dX_{r}+\varepsilon^{\tilde{\mu}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\int_{s}^{t}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\sigma(X_{s})+\sigma^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\sigma^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dX_{r}+\varepsilon^{\sigma}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{u}^{\mathbb{P}}.$
@@
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{u}^{\mathbb{P}}$
$\qquad+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{u}^{\mathbb{P}}$
@@
$=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mu}(X_{r})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}drdu+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma(X_{r})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{r}^{\mathbb{P}}du$
$\qquad+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mu}(X_{r})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dr\circ dB_{u}^{\mathbb{P}}+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sigma(X_{r})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$\qquad+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{u}^{\mathbb{P}}$
@@
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du$
$\qquad+(\tilde{\mu}^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dX_{v}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dX_{v}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{r}^{\mathbb{P}}du$
$\qquad+\sigma^{\prime}(X_{s})\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dX_{v}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dr\circ dB_{u}^{\mathbb{P}}+(\sigma^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\circ dX_{v}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$\qquad+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{u}^{\mathbb{P}}$
@@
$=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du$
$\qquad+(\tilde{\mu}^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}drdu$
$\qquad+\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{r}^{\mathbb{P}}du$
$\qquad+\sigma^{\prime}(X_{s})\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dr\circ dB_{u}^{\mathbb{P}}$
$\qquad+(\sigma^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$\qquad+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}du$
$\qquad+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ dB_{u}^{\mathbb{P}}$
@@
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du$

### v2

@@ Appendix A Proof of Step 2
$+\psi_{21}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ dB^{\mathbb{Q}}_{r}du+\psi_{22}^{s,t}\int_{s}^{t}\int_{s}^{u}\circ dB^{\mathbb{Q}}_{r}\circ dB_{u}^{\mathbb{Q}}$
$+\psi_{222}^{s,t}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dB_{v}^{\mathbb{Q}}\circ dB_{r}^{\mathbb{Q}}\circ dB_{u}^{\mathbb{Q}}+\mathsf{\varepsilon}^{\Psi_{2}},$
@@
Appendix B Proof of Step 3
Applying stochastic Taylor expansion on $\tilde{\mu}$ and $\sigma$ around $X_{s}$, and truncating the Taylor expansion at level 2, we have the following approximation of $\tilde{\mu}$ and $\sigma$,
$\tilde{\mu}(X_{u})$ $=\tilde{\mu}(X_{s})+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dX_{r}$ (37)
$+\underbrace{\int_{s}^{u}\int_{s}^{r}\int_{s}^{v}\tilde{\mu}^{\prime\prime\prime}(\xi_{\tau})\circ dX_{\tau}\circ dX_{v}\circ dX_{r}}_{\varepsilon^{\tilde{\mu}}}$
$\sigma(X_{u})$ $=\sigma(X_{s})+\sigma^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\sigma^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dX_{r}$ (38)
$+\underbrace{\int_{s}^{u}\int_{s}^{r}\int_{s}^{v}\sigma^{\prime\prime\prime}(\eta_{\tau})\circ dX_{\tau}\circ dX_{v}\circ dX_{r}}_{\varepsilon^{\sigma}}$
@@ Appendix B Proof of Step 3
$X_{t}$ $=X_{s}+\int_{s}^{t}\left(\tilde{\mu}(X_{s})+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}{\circ}dX_{v}\circ dX_{r}+\varepsilon^{\tilde{\mu}}\right)du$
$+\int_{s}^{t}\left(\sigma(X_{s})+\sigma^{\prime}(X_{s})\int_{s}^{u}\circ dX_{r}+\sigma^{\prime\prime}(X_{s})\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dX_{r}+\varepsilon^{\sigma}\right)\circ dB_{u}^{\mathbb{P}}.$
@@
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}$
$\qquad+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)du$
$\qquad+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)du$
$\qquad+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\circ dB_{u}^{\mathbb{P}}$
$\qquad+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\circ dB_{u}^{\mathbb{P}}$
@@
$=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\tilde{\mu}(X_{r})drdu$
$\qquad+\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}du+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\tilde{\mu}(X_{r})dr\circ dB_{u}^{\mathbb{P}}$
$\qquad+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\tilde{\mu}(X_{r})dr\circ dB_{u}^{\mathbb{P}}+\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$\qquad+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)du$
$\qquad+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\circ dB_{u}^{\mathbb{P}}$
@@
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}$
$+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du$
$+\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}dr\circ dB_{u}^{\mathbb{P}}+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$+(\tilde{\mu}^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dX_{v}drdu+\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dB_{r}^{\mathbb{P}}du$
$+\sigma^{\prime}(X_{s})\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dX_{v}dr\circ dB_{u}^{\mathbb{P}}$
$+(\sigma^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\circ dX_{v}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)du$
$+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\circ dB_{u}^{\mathbb{P}}$
@@
$=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu$
$+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du+\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}dr\circ dB_{u}^{\mathbb{P}}$
$+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$+(\tilde{\mu}^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)drdu$
$+\tilde{\mu}^{\prime}(X_{s})\sigma^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\circ dB_{r}^{\mathbb{P}}du$
$+\sigma^{\prime}(X_{s})\tilde{\mu}^{\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)dr\circ dB_{u}^{\mathbb{P}}$
$+(\sigma^{\prime}(X_{s}))^{2}\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$
$+\tilde{\mu}^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)du$
$+\sigma^{\prime\prime}(X_{s})\int_{s}^{t}\int_{s}^{u}\int_{s}^{r}\left(\tilde{\mu}(X_{v})dv+\sigma(X_{v})\circ dB_{v}^{\mathbb{P}}\right)\left(\tilde{\mu}(X_{r})dr+\sigma(X_{r})\circ dB_{r}^{\mathbb{P}}\right)\circ dB_{u}^{\mathbb{P}}$
$+\varepsilon^{\tilde{\mu}}\int_{s}^{t}du+\varepsilon^{\sigma}\int_{s}^{t}\circ dB^{\mathbb{P}}_{u}$
@@
$X_{t}$ $=X_{s}+\tilde{\mu}(X_{s})\int_{s}^{t}du+\sigma(X_{s})\int_{s}^{t}\circ dB_{u}^{\mathbb{P}}+\tilde{\mu}^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}drdu$
$\qquad+\tilde{\mu}^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}du+\sigma^{\prime}(X_{s})\tilde{\mu}(X_{s})\int_{s}^{t}\int_{s}^{u}dr\circ dB_{u}^{\mathbb{P}}$
$\qquad+\sigma^{\prime}(X_{s})\sigma(X_{s})\int_{s}^{t}\int_{s}^{u}\circ dB_{r}^{\mathbb{P}}\circ dB_{u}^{\mathbb{P}}$